Find the molality of a solution made by mixing 85.3 g of 1.35 M HCl, 0.867 g of HCl and 89.1 g of
water.
1
Expert's answer
2021-09-22T00:27:17-0400

Molar mass of HCl = 36.458 g/mol


Moles of HCl"=\frac{85.3}{36.458}=2.34 moles"


Volume of "HCl=\frac{2.34}{1.35}=1.73m^3"


Moles of water"=\frac{89.1}{18.02}=4.94 moles"



Molarity =moles/Volume


Molarity "=\frac{4.94}{1.73}=2.86M"